OpenStudy (anonymous):

how do you find ∫ -3 to 3 of sqrt(9-(x^2))?

OpenStudy (anonymous):

i know you have to find the antiderivative first but like how do you do that for this?

OpenStudy (freckles):

\[y=\sqrt{9-x^2} \\ y^2=9-x^2 \\ x^2+y^2=9 \\ \\] x^2+y^2=9 is a circle with center (0,0) and radius 3 y=sqrt(9-x^2) is the top part of the circle with center (0,0) and radius 3 find the area enclosed by the circle x^2+y^2=9 and half it since you have half a circle that is you have this:  pretend I drew a lovely half circle for you